OpenStudy (anonymous):

Which graph represents the solution set for the inequality 16 ≥ -3x + 4 > -17?

OpenStudy (amoodarya):

16 ≥ -3x + 4 > -1 first multiply by -1 careful about changing the sign -16 <= +3x- 4 <+1

OpenStudy (amoodarya):

now plus 4 -16 +4<= +3x- 4+4 <+1+4 -12 <=3x<5

OpenStudy (amoodarya):

and now divide by 3

OpenStudy (amoodarya):

-12 <=3x<5 --->/3 -4<=x<5/3